OpenStudy (melissa_something):

4^x+8=7: Solve & express in terms of natural logarithms. WILL MEDAL!

OpenStudy (anonymous):

subtract 8 from both sides, and take a natural log. you get no solution

OpenStudy (anonymous):

ALSO, for 4^x+8=7 to be true, 4^x must =-1 and no exponent will make a pos. number to a negative number.

Nnesha (nnesha):

is that \[\huge\rm 4^{x+8} = 7\] ??

OpenStudy (anonymous):

that could also be it:)

OpenStudy (melissa_something):

Uhh ok lemme work this out real quick. & Yes @Nnesha it is!

Nnesha (nnesha):

yes so you can change exponential to log form